OpenStudy (anonymous):

Find the 37th term of the following sequence. a1= -3 ; d = 2.8

OpenStudy (anonymous):

N=(l-a/d)+1

OpenStudy (anonymous):

Term n =(a+(N-1)d)

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{37} = -3 + (37 - 1)2.8\]

OpenStudy (anonymous):

\[a _{37} = -3 + (36)2.8\]

OpenStudy (anonymous):

@some_someone. Why can't you just leave after telling the formula.

OpenStudy (anonymous):

why does it matter^

OpenStudy (anonymous):

@sathishSON well that was rude.

OpenStudy (anonymous):

U have to try yourself with the given formula....

OpenStudy (anonymous):

dude nobody on here is trying to learn. we want answers so stop with the rude comments and get a life

OpenStudy (anonymous):

Ok....cool

OpenStudy (anonymous):

thanks for helping, but don't be rude. u look dumb af

OpenStudy (anonymous):

Ok....guys.

OpenStudy (anonymous):

ACTUALLY WE ARE HERE TO HELP NOT GIVE ANSWERS!

OpenStudy (anonymous):

HELP AND GUIDE USERS TO THE ANSWERS, NOT JUST GIVE THE ANSWER(S)!

OpenStudy (anonymous):

That's good !!!

OpenStudy (anonymous):

INDEED! ;D

OpenStudy (anonymous):

Ok....I'm jumping to history.

OpenStudy (anonymous):

OKAII BAII :)